OpenStudy (anonymous):

In triangle ABC, what is the length of side b if B = 10°, C = 135°, and c = 45? 3.24 3.33 3.82 11.05

OpenStudy (anonymous):

apply law of sins: sinC/c = sinB/b

OpenStudy (anonymous):

I set it up like this and got http://www.wolframalpha.com/input/?i=sin^-1%28sin135%2810%29%2F%2845

OpenStudy (anonymous):

Yeah, looks the right way.

OpenStudy (anonymous):

it looks the right way but..

OpenStudy (anonymous):

You got me confused too.\[b = \frac {\sin{\beta} \times c}{\sin{\gamma}}\]

OpenStudy (anonymous):

you know what I have no idea no more

OpenStudy (anonymous):

The law says that the ratio between lenght and sin of the opposed angle is the same for all three sides / angles.

OpenStudy (anonymous):

im brain dead now